# Bispectrum Islands

### v1

@@ I Introduction
This work aims to chart the landscape of all possible bispectra induced by interactions with a hidden sector during inflation, consisting of degrees of freedom beyond the inflaton and graviton. Our approach is model-independent and, much like the S-matrix positivity bound/bootstrap programs de Rham et al. (2017, 2018, 2019); Paulos et al. (2019); Arkani-Hamed et al. (2021); Tolley et al. (2021); Caron-Huot and Van Duong (2021); Bellazzini et al. (2021); Caron-Huot et al. (2021); Guerrieri and Sever (2021); de Rham et al. (2022); Kruczenski et al. (2022), it is grounded in fundamental principles such as unitarity, locality, analyticity and symmetry. [^2] See Refs. de Rham and Melville (2017); Grall and Melville (2020, 2022); Grall (2023) for constraints on cosmological EFTs based on unitarity and Grall et al. (2020a, b) for symmetry ones. These principles are used to directly constrain the observable—the bispectrum—without making explicit reference to the hidden sector’s Lagrangian.
@@
While in this work we do not claim to derive the most general bounds on the bispectrum, our goal is to demonstrate that, by leveraging the consequences of unitarity and isometries in rigid de Sitter space, it is in principle possible to impose nontrivial constraints directly on inflationary correlators—much like how S-matrices are constrained in flat space using the same principles in Minkowski space. See also Hogervorst et al. (2023); Di Pietro et al. (2021, 2023); Penedones et al. (2023); Loparco et al. (2023, 2025), where the same principles are employed to constrain quantum field theories in de Sitter.
@@ II General framework
×$\times\left(e_{1}^{5}-3e_{1}^{3}e_{2}-4e_{1}e_{2}^{2}+11e_{1}^{2}e_{3}-4e_{2}e_{3}\right)\,.$
@@
where $\alpha$ is a dimensionless coupling and $\Lambda$ is the cut-off scale. As noted earlier, aside from a few assumptions, we treat ${\cal O}$ as a fully generic scalar operator. The assumptions underlying our bounds are as follows: we require that the correlators of ${\cal O}$ satisfy the Ward identities associated with the de Sitter isometry group $\text{SO}(4,1)$. This is ensured by demanding that the hidden sector is not strongly coupled to the time-dependent background, and can be treated as a spectator operator living in a fixed de Sitter spacetime. In particular, we demand that, in the absence of the weak coupling above, the vacuum two-point function of ${\cal O}$ is given by the Källén–Lehmann (KL) spectral decomposition of a scalar operator in $\text{dS}_{4}$ (see e.g. Loparco et al. (2023)), which takes the form [^3] Note that this follows from the assumption that the expectation value $\langle{\cal O}(\eta,\bm{x})\rangle$ vanishes. Symmetries require any nonzero $\langle{\cal O}\rangle$ to be constant, so without loss of generality, one can always redefine ${\cal O}$ to make this true.
@@
Here, $\mu\equiv\left(\frac{m^{2}}{H^{2}}-\frac{9}{4}\right)^{1/2}$ is the mass-index, $\sigma$ is the geodesic distance between the bulk points $(\eta,\bm{x})$ and $(\eta^{\prime},\bm{x}^{\prime})$, namely
@@
×$\times\,_{2}F_{1}\left(\frac{3}{2}+i\mu;\frac{3}{2}-i\mu;2,\dfrac{1+\sigma}{2}\right)\,,$ (14)
@@
A brief comment on the exclusion of light fields is in order. Since our goal is to constrain scenarios where a single scalar degree of freedom sources the adiabatic perturbations, we must ensure that $\delta\phi$ is the only massless mode in the theory. This requires a gap in the hidden sector spectrum such that $\rho(m)=0$ for $m<m_{\text{gap}}$. However, imposing a sharp gap above the threshold $23H/$ 2 is incompatible with the spectrum of generic interacting QFTs in de Sitter, which at minimum spans the entire principal series. That interactions generically render the spectrum gapless within the principal series is evident already at the perturbative level. A simple example is the one-loop correction to the two-point function of an elementary heavy field in de Sitter, which induces a non-zero spectral density even below the particle production threshold Di Pietro et al. (2021); Hogervorst et al. (2023). While in principle a sharp gap can be imposed within the complementary series, i.e. $m_{\text{gap}}<3H/2$, this comes at the cost of introducing an arbitrary scale $m_{\text{gap}}$ into the resulting positivity bounds. For this reason, we choose to impose $m_{\text{gap}}=3H/2$. This is a natural gap in de Sitter, in the sense that as long as the two-point function of ${\cal O}$ decays faster than $|\sigma|^{3/2}$ at large spatial separations ($\sigma\to-\infty$), the principal series provides a complete basis for the spectral decomposition Loparco et al. (2023). Furthermore, we expect this gap to be stable under quantum corrections.
@@
The leading-order diagrams contributing to the bispectrum are depicted in Figure 1. They include an exchange process of order ${\cal O}(\dot{\bar{\phi}})$ induced by the two-point function of the operator ${\cal O}$ and another diagram of order ${\cal O}(\dot{\bar{\phi}}^{3})$, proportional to the three-point function $\langle{\cal O}^{3}\rangle$. Our next primary ingredient for deriving the bounds on the bispectrum is that the former contribution always dominates. Details about the typical size of each contribution will follow in Section IV. Our theoretical reason for excluding the three-vertex diagram from our analysis is that, unlike the two-point function of the hidden sector, there is no useful dispersion relation to similarly constrain the analytic structure of the three-point function $\langle\mathcal{O}(x)\mathcal{O}(y)\mathcal{O}(z)\rangle$. To make progress in constraining the final bispectrum, we therefore focus on setups in which the contribution from the three-vertex diagram is subdominant, and the bispectrum can be expressed entirely in terms of the two-point function of $\mathcal{O}$, whose form is fully fixed by the Källén–Lehmann representation.
@@ III Bispectrum from Operator Exchange
We employ the in-in formalism to compute the two-vertex exchange diagram shown in Figure 1. The key distinction from standard in-in calculations lies in the treatment of the hidden sector: within the interaction picture, the Wick theorem does not apply to the operator ${\cal O}$. In other words, ${\cal O}$ should not be treated as a free scalar field with a fixed mass, unlike factors of $\delta\phi$, which can be contracted using the usual rules. Instead, at each order in perturbation theory, one must insert the appropriate Green functions for ${\cal O}$. These Green functions are generally unknown unless the full hidden-sector theory is specified, but they are invariant under the action of the de Sitter group.
@@
$G^{\cal O}_{+-}(\eta,\eta^{\prime},\bm{x}-\bm{x}^{\prime})$ $=\langle 0|\,{\cal O}(\eta,\bm{x}){\cal O}(\eta^{\prime},\bm{x}^{\prime})|0\rangle\,,$
@@
where $B(k_{1},k_{2},k_{3};\mu)$ is the bispectrum induced by the exchange of a heavy scalar $f$ with the mass-index $\mu$, which is coupled to $\delta\phi$ in the same way as the original scalar operator ${\cal O}$ (given in Eq. (15)). In conclusion, with a known spectral density $\rho(\mu)$, the above representation reduces the computation of the full bispectrum to evaluating the single-particle exchange contributions $B(k_{i};\mu)$.
@@
In recent years, powerful techniques from the cosmological bootstrap have yielded useful analytical expressions for $B(k_{1},k_{2},k_{3};\mu)$ Baumann et al. (2024); Arkani-Hamed et al. (2020). In particular, it has been shown that the bispectrum exchange diagram can be easily mapped to the exchange diagram of a four-point function involving an auxiliary, conformally coupled field $\varphi$, which interacts with the heavy scalar field $f$ through a cubic term of the form $\varphi^{2}\,f$. Here we only quote the final formulae for this seed four-point function, which exhibits a much simpler analytic structure compared to the original three-point diagram. The bispectrum will then simply follow by acting with an appropriate weight-shifting operator Baumann et al. (2020).
@@
in which $\eta_{0}$ is the (conformal) time at which the correlator is evaluated, $s=|\bm{k}_{1}+\bm{k}_{2}|$ is the s-channel exchanged momentum, and ${\cal F}$ is given by
@@
×$\times\int_{-\infty}^{0}\frac{\mathrm{d}\eta}{\eta^{2}}e^{\pm ik_{12}\eta}\int_{-\infty}^{0}\frac{\mathrm{d}\eta^{\prime}}{\eta^{\prime 2}}e^{\pm ik_{34}\eta^{\prime}}G_{\pm\pm}(s,\eta,\eta^{\prime})\,.$
@@
×$\times\left((\beta_{0}-1)\alpha_{+}\,F_{-}(u)-\alpha_{-}(\beta_{0}+1)\hat{F}_{+}(u)\right)\,,$ (29)
@@
$\lim_{k_{3}\ll k1\sim k_{2}}B_{s}(k_{1},k_{2},k_{3};\mu)=\kappa\dfrac{1}{k_{1}^{6}}\left(\dfrac{k_{3}}{k_{1}}\right)^{-3/2+i\mu}$
×$\,\,\,\times\beta(\mu)(-15+4\mu(-4i+\mu))2^{-7/2-i\mu}+{\rm c.c.},$ (31)
@@
[…]
×$\times\Gamma(-i\mu)\Gamma(2i\mu)(-i+\sinh(\pi\mu))\,.$ (32)
@@ IV The bispectrum sign and size
which must be negative. This property follows directly from the fact that the bispectra associated with the exchange of individual states are negative for all masses (as can be seen in Figure 2), combined with the positivity of the spectral density, $\rho(\mu)>0$. As we will soon see, the bispectrum $B(k_{1},k_{2},k_{3})$ can become positive for sufficiently squeezed triangles, with its sign oscillating due to the superposition of cosmological collider oscillations (see Eq. (31)) induced by light fields in the spectrum, i.e. $\mu\lesssim 1$.
@@
Finally, we revisit our assumption that the diagram proportional to $\langle{\cal O}^{3}\rangle$ in Figure 1 is sub-dominant, relative to the single-exchange process. The corresponding amplitude of non-Gaussianity is approximately given by
@@ V Bounds on the bispectrum shape
To derive the optimal values of ${\cal S}$ for a given configuration $(k_{2}/k_{1},k_{3}/k_{1})$, we first note that $\rho(\mu)$ can always be rescaled so that the denominator is fixed to $-1$. In other words, since the shape is scale invariant, without loss of generality, we can always rescale the density to impose the following constraint on $\rho(\mu)$:
@@
Even with this simple final ansatz (48), the complexity of the bispectrum templates $B(k_{i};\mu)$ prevents us from obtaining analytical expressions for the optimal values of the shape function. Instead, for each configuration $(k_{2}/k_{1},k_{3}/k_{1})$, we numerically evaluate the above ansatz and extract the optimal masses $\mu_{\text{max}}$ and $\mu_{\text{min}}$, along with the corresponding extrema of the shape function, ${\cal S}_{\text{max}}$ and ${\cal S}_{\text{min}}$—see Figure 3.
In Figure 4, we have presented our final bounds ${\cal S}_{\text{max,min}}$ for isosceles triangles (with $k_{2}=k_{3}$), as a function of the squeezing parameter $k_{3}/k_{1}$. Let us summarize the general properties of our bounds:
  • We have numerically verified that for all triangle configurations the upper bound on ${\cal S}$ is always saturated by the bispectrum shape corresponding to the exchange of a heavy state with mass $m=3H/2$ (i.e. $\mu=0$), see Figure 3. In other words, we find $\mu_{\text{max}}=0$ and:
@@
  • In contrast to $\mu_{\text{max}}$, $\mu_{\text{min}}$ depends on the ratio $k_{3}/k_{1}$, as can be inferred from Figure 3. In other words, the lower curve in Figure 4, ${\cal S}_{\text{min}}$, does not correspond to the bispectrum shape resulting from the exchange of any mass eigenstate. We also observe that the value of $\mu_{\text{min}}$ ranges from 0 (in the squeezed limit) up to $\mu\approx 3.0$, near the folded configuration.
  • Moreover, we find that the lower bound on the bispectrum shape remains positive as long as $k_{3}/k_{1}\gtrsim 0.027$, but becomes negative for more squeezed triangles. Combined with the earlier result that the equilateral bispectrum is negative (i.e. $B(1,1,1)<0$), this implies that the overall sign of the bispectrum is negative for a range of triangle configurations—as long as they are not too squeezed. The corresponding threshold depends on the angle between the short and the long mode, see Figure 5.
Figure 3: The ansatz for the bispectrum shape ${\cal S}(k_{i};\mu)$ is shown as a function of $\mu$, defined by Eq. (48). As the plots illustrate across various configurations, the upper bound on the bispectrum shape is always set by the ansatz evaluated at $\mu=0$, while the lower bound ${\cal S}_{\text{min}}$ corresponds to the exchange of fields with masses varying in the range $30<\mu_{\text{min}}<$ 3.

### v2

@@ I Introduction
This work aims to chart the landscape of all possible bispectra induced by interactions with a hidden sector during inflation, consisting of degrees of freedom beyond the inflaton and graviton. Our approach is model-independent and, much like the S-matrix positivity bound/bootstrap programs de Rham et al. (2017, 2018, 2019); Paulos et al. (2019); Arkani-Hamed et al. (2021); Tolley et al. (2021); Caron-Huot and Van Duong (2021); Bellazzini et al. (2021); Caron-Huot et al. (2021); Guerrieri and Sever (2021); de Rham et al. (2022); Kruczenski et al. (2022), it is grounded in fundamental principles such as unitarity, locality, analyticity and symmetry. [^2] See Refs. de Rham and Melville (2017); Grall and Melville (2020, 2022); Grall (2023); Stefanyszyn et al. (2024) for constraints on cosmological EFTs based on unitarity, Grall et al. (2020a, b); Agui Salcedo and Melville (2023) for symmetry ones and Melville and Pimentel (2024a, b); Cespedes and Jazayeri (2025) for progress in relating cosmological correlators with their flat counterparts. These principles are used to directly constrain the observable—the bispectrum—without making explicit reference to the hidden sector’s Lagrangian.
@@
While in this work we do not claim to derive the most general bounds on the bispectrum, our goal is to demonstrate that, by leveraging the consequences of unitarity and isometries in rigid de Sitter space, it is in principle possible to impose nontrivial constraints directly on inflationary correlators—much like how S-matrices are constrained in flat space using the same principles in Minkowski space. See also Hogervorst et al. (2023); Di Pietro et al. (2021, 2023); Penedones et al. (2023); Loparco et al. (2023, 2025), where the same principles are employed to constrain quantum field theories in de Sitter space.
@@ II General framework
$\times\left(e_{1}^{5}-3e_{1}^{3}e_{2}-4e_{1}e_{2}^{2}+11e_{1}^{2}e_{3}-4e_{2}e_{3}\right)\,.$
@@
where $\alpha$ is a dimensionless coupling and $\Lambda$ is the cutoff scale. As noted earlier, aside from a few assumptions, we treat $\cal O$} as a fully generic scalar operator. The assumptions underlying our bounds are as follows: we require that the correlators of $\cal O$} satisfy the Ward identities associated with the de Sitter isometry group $\text{SO}(4,1)$. This is ensured by demanding that the hidden sector is not strongly coupled to the time-dependent background, and can be treated as a spectator operator living in a fixed de Sitter spacetime. In particular, we demand that, in the absence of the weak coupling above, the vacuum two-point function of $\cal O$} is given by the Källén–Lehmann (KL) spectral decomposition of a scalar operator in $\text{dS}_{4}$ (see e.g. Loparco et al. (2023)), which takes the form [^3] Note that this follows from the assumption that the expectation value $\langle{\cal O}(\eta,\bm{x})\rangle$ vanishes. Symmetries require any nonzero $\langle{\cal O}\rangle$ to be constant, so without loss of generality, one can always redefine $\cal O$} to make this true.
@@
Here, $\mu\equiv\left(\frac{m^{2}}{H^{2}}-\frac{9}{4}\right)^{1/2}$ is the mass index, $\sigma$ is the geodesic distance between the bulk points $(\eta,\bm{x})$ and $(\eta^{\prime},\bm{x}^{\prime})$, namely
@@
$\times\,_{2}F_{1}\left(\frac{3}{2}+i\mu;\frac{3}{2}-i\mu;2,\dfrac{1+\sigma}{2}\right)\,$, (14)
@@
A brief comment on the exclusion of light fields is in order. Since our goal is to constrain scenarios where a single scalar degree of freedom sources the adiabatic perturbations, we must ensure that $\delta\phi$ is the only massless mode in the theory. This requires a gap in the hidden sector spectrum such that $\rho(m)=0$ for $m<m_{\text{gap}}$. However, imposing a sharp gap above the threshold $3H/2$ is incompatible with the spectrum of generic interacting QFTs in de Sitter, which at minimum spans the entire principal series. That interactions generically render the spectrum gapless within the principal series is evident already at the perturbative level. A simple example is the one-loop correction to the two-point function of an elementary heavy field in de Sitter, which induces a non-zero spectral density even below the particle production threshold Di Pietro et al. (2021); Hogervorst et al. (2023). While in principle a sharp gap can be imposed within the complementary series, i.e. $m_{\text{gap}}<3H/2$, this comes at the cost of introducing an arbitrary scale $m_{\text{gap}}$ into the resulting positivity bounds. For this reason, we choose to impose $m_{\text{gap}}=3H/2$. This is a natural gap in de Sitter, in the sense that as long as the two-point function of $\cal O$} decays faster than $|\sigma|^{3/2}$ at large spatial separations ($\sigma\to$-$\infty$), the principal series provides a complete basis for the spectral decomposition Loparco et al. (2023). Furthermore, we expect this gap to be stable under quantum corrections.
@@
The leading-order diagrams contributing to the bispectrum are depicted in Fig. 1. They include an exchange process of order ${\cal O}(\dot{\bar{\phi}})$ induced by the two-point function of the operator $\cal O$} and another diagram of order ${\cal O}(\dot{\bar{\phi}}^{3})$, proportional to the three-point function $\langle{\cal O}^{3}\rangle$. Our next primary ingredient for deriving the bounds on the bispectrum is that the former contribution always dominates. Details about the typical size of each contribution will follow in Section IV. Our theoretical reason for excluding the three-vertex diagram from our analysis is that, unlike the two-point function of the hidden sector, there is no useful dispersion relation to similarly constrain the analytic structure of the three-point function $\langle\mathcal{O}(x)\mathcal{O}(y)\mathcal{O}(z)\rangle$. To make progress in constraining the final bispectrum, we therefore focus on setups in which the contribution from the three-vertex diagram is subdominant, and the bispectrum can be expressed entirely in terms of the two-point function of $\mathcal{O}$, whose form is fully fixed by the Källén–Lehmann representation.
@@ III Bispectrum from Operator Exchange
We employ the in-in formalism to compute the two-vertex exchange diagram shown in Fig. 1. The key distinction from standard in-in calculations lies in the treatment of the hidden sector: within the interaction picture, the Wick theorem does not apply to the operator $\cal O$}. In other words, $\cal O$} should not be treated as a free scalar field with a fixed mass, unlike factors of $\delta\phi$, which can be contracted using the usual rules. Instead, at each order in perturbation theory, one must insert the appropriate Green functions for $\cal O$}. These Green functions are generally unknown unless the full hidden-sector theory is specified, but they are invariant under the action of the de Sitter group.
@@
$G^{\cal O}_{-+}(\eta,\eta^{\prime},\bm{x}-\bm{x}^{\prime})$ $=\langle 0|\,{\cal O}(\eta,\bm{x}){\cal O}(\eta^{\prime},\bm{x}^{\prime})|0\rangle\,$,
@@
where $B(k_{1},k_{2},k_{3};\mu)$ is the bispectrum induced by the exchange of a heavy scalar $f$ with the mass index $\mu$, which is coupled to $\delta\phi$ in the same way as the original scalar operator $\cal O$} (given in Eq. (15)). In conclusion, with a known spectral density $\rho(\mu)$, the above representation reduces the computation of the full bispectrum to evaluating the single-particle exchange contributions $B(k_{i};\mu)$.
@@
In recent years, powerful techniques from the cosmological bootstrap have yielded useful analytical expressions for $B(k_{1},k_{2},k_{3};\mu)$ Baumann et al. (2024); Arkani-Hamed et al. (2020). In particular, it has been shown that the bispectrum exchange diagram can be easily mapped to the exchange diagram of a four-point function involving an auxiliary, conformally coupled field $\varphi$, which interacts with the heavy scalar field $f$ through a cubic term of the form $\varphi^{2}\,f$. Here we only quote the final formulas for this seed four-point function, which exhibits a much simpler analytic structure compared to the original three-point diagram. The bispectrum will then simply follow by acting with an appropriate weight-shifting operator Baumann et al. (2020).
@@
[…]
@@
$\times\int_{-\infty}^{0}\frac{\mathrm{d}\eta}{\eta^{2}}e^{\pm ik_{12}\eta}\int_{-\infty}^{0}\frac{\mathrm{d}\eta^{\prime}}{\eta^{\prime 2}}e^{\pm ik_{34}\eta^{\prime}}G_{\pm\pm}(s,\eta,\eta^{\prime})\,.$
@@
$\times\left((\beta_{0}-1)\alpha_{+}\,F_{-}(u)-\alpha_{-}(\beta_{0}+1)\hat{F}_{+}(u)\right)\,$, (29)
@@
$\lim_{k_{3}\ll k1\sim k_{2}}B(k_{1},k_{2},k_{3};\mu)=\kappa\dfrac{1}{k_{1}^{6}}\left(\dfrac{k_{3}}{k_{1}}\right)^{-3/2+i\mu}$
$\,\,\,\times\beta(\mu)(-15+4\mu(-4i+\mu))2^{-7/2-i\mu}+{\rm c.c.},$ (31)
@@
$\beta(\mu)$ $=-\dfrac{\pi\mu\,2^{1/2-3i\mu}}{2\cosh(\pi\mu)\Gamma(1+i\mu)}$
$\times\Gamma(-i\mu)\Gamma(2i\mu)(-i+\sinh(\pi\mu))\,.$ (32)
@@ IV The bispectrum sign and size
which must be negative. This property follows directly from the fact that the bispectra associated with the exchange of individual states are negative for all masses (as can be seen in Fig. 2), combined with the positivity of the spectral density, $\rho(\mu)>0$. As we will soon see, the bispectrum $B(k_{1},k_{2},k_{3})$ can become positive for sufficiently squeezed triangles, with its sign oscillating due to the superposition of cosmological collider oscillations (see Eq. (31)) induced by light fields in the spectrum, i.e. $\mu\lesssim 1$.
@@
Finally, we revisit our assumption that the diagram proportional to $\langle{\cal O}^{3}\rangle$ in Fig. 1 is sub-dominant, relative to the single-exchange process. The corresponding amplitude of non-Gaussianity is approximately given by
@@ V Bounds on the bispectrum shape
To derive the optimal values of $\cal S$} for a given configuration $(k_{2}/k_{1},k_{3}/k_{1})$, we first note that $\rho(\mu)$ can always be rescaled so that the denominator is fixed to 1-1. In other words, since the shape is scale invariant, without loss of generality, we can always rescale the density to impose the following constraint on $\rho(\mu)$:
@@
Even with this simple final ansatz (48), the complexity of the bispectrum templates $B(k_{i};\mu)$ prevents us from obtaining analytical expressions for the optimal values of the shape function. Instead, for each configuration $(k_{2}/k_{1},k_{3}/k_{1})$, we numerically evaluate the above ansatz and extract the optimal masses $\mu_{\text{max}}$ and $\mu_{\text{min}}$, along with the corresponding extrema of the shape function, ${\cal S}_{\text{max}}$ and ${\cal S}_{\text{min}}$—see Fig. 3.
In Fig. 4, we have presented our final bounds ${\cal S}_{\text{max,min}}$ for isosceles triangles (with $k_{2}=k_{3}$), as a function of the squeezing parameter $k_{3}/k_{1}$. Let us summarize the general properties of our bounds:
  • We have numerically verified that for all triangle configurations the upper bound on $\cal S$} is always saturated by the bispectrum shape corresponding to the exchange of a heavy state with mass $m=3H/2$ (i.e. $\mu=0$), see Fig. 3. In other words, we find $\mu_{\text{max}}=0$ and:
@@
  • In contrast to $\mu_{\text{max}}$, $\mu_{\text{min}}$ depends on the ratio $k_{3}/k_{1}$, as can be inferred from Fig. 3. In other words, the lower curve in Fig. 4, ${\cal S}_{\text{min}}$, does not correspond to the bispectrum shape resulting from the exchange of any mass eigenstate. We also observe that the value of $\mu_{\text{min}}$ ranges from 0 (in the squeezed limit) up to $\mu\approx 3.0$, near the folded configuration.
  • Moreover, we find that the lower bound on the bispectrum shape remains positive as long as $k_{3}/k_{1}\gtrsim 0.027$, but becomes negative for more squeezed triangles. Combined with the earlier result that the equilateral bispectrum is negative (i.e. $B(1,1,1)<0$), this implies that the overall sign of the bispectrum is negative for a range of triangle configurations—as long as they are not too squeezed. The corresponding threshold depends on the angle between the short and the long mode, see Fig. 5.
Figure 3: The ansatz for the bispectrum shape ${\cal S}(k_{i};\mu)$ is shown as a function of $\mu$, defined by Eq. (48). As the plots illustrate across various configurations, the upper bound on the bispectrum shape is always set by the ansatz evaluated at $\mu=0$, while the lower bound ${\cal S}_{\text{min}}$ corresponds to the exchange of fields with masses varying in the range $0<\mu_{\text{min}}<3$.